6. Let W be the set of all vectors of the form W {(a,b,c): a – 2b + 4z = 0} Is W a subspace of the vector space V = R3?
Answer #1

Hey there stay home and stay safe

\small W=\{(a,b,c)|a-2b+4c=0\}\\ $Clearly seen $(0,0,0)\in W, $ W is non-empty $\\ Let, \alpha=(a_1,b_1,c_1),\beta=(a_2,b_2,c_2)\in W\\ Then,a_1-2b_1+4c_1=0,a_2-2b_2+4c_2=0\\ \alpha+\beta=(a_1+a_2,b_1+b_2,c_1+c_2)\in W\\ Since, (a_1+a_2)-2(b_1+b_2)+4(c_1+c_2)=a_1-2b_1+4c_1+a_2-2b_2+4c_2=0+0=0\\ Let,c\in R(scalar), c\alpha=c(a_1,b_1,c_1)=(ca_1,cb_1,cc_1)\in W\\ Since, ca_1-2cb_1+4cc_1=c(a_1-2b_1+4c_1)=c(0)=0\\ $Therefore W is a subspace$\\

And please leave a like. Thank you.